OpenStudy (lgbasallote):

Find the center and radius of the circle with equation x^2 + y^2 - 6x + 10y + 9 = 0

OpenStudy (raden):

if given a circle equation : x^2+y^2+AX+By+C=0, then the centre of circle at point : (a,b) with a = -A/2 and b = -B/2 for radius (r) = sqrt((1/4)A^2 + (1/4)B^2 - C)

OpenStudy (lgbasallote):

^?

OpenStudy (anonymous):

I like the explanation on this page: http://www.analyzemath.com/Calculators/Cen_Rad_Cal.html

OpenStudy (lgbasallote):

...algebra.....scary......

OpenStudy (lgbasallote):

too many variables

OpenStudy (raden):

A=-6, B=10, C=9

OpenStudy (anonymous):

Basically, you can look at the \(x^2 + ax\) part of the equation and use completing the square to rewrite it. Then you can look at the \(y^2 +by\) part of the equation and use completing the square to rewrite that. If you don't love completing the square, just use the formula given. Gather all of the constant terms on one side of the equation, and suddenly the equation is in standard form.

OpenStudy (lgbasallote):

completing squares huh..

OpenStudy (anonymous):

\(x^2 + a x = (x + a/2)^2 - (a/2)^2 \) For your problem, a=-6

OpenStudy (lgbasallote):

hmm

OpenStudy (anonymous):

How familiar are you with completing the square? I'm sure you've seen it come up.

OpenStudy (lgbasallote):

not that good

OpenStudy (raden):

i mean, A,B, and C are constanta

OpenStudy (lgbasallote):

i'll try my best though \[x^2 - 6x + y^2 + 10y = -9\] \[(x - 3)^2 + (y+ 5)^2 = -9 + 9 + 25\]

OpenStudy (lgbasallote):

so then it becomes \[(x-3)^2 + (y+5)^2 = 25\]

OpenStudy (raden):

it will work also

OpenStudy (lgbasallote):

so then the center is (3,-5) and radius is 5...

OpenStudy (raden):

yes, u are right

OpenStudy (lgbasallote):

feel free to call me wrong though

OpenStudy (lgbasallote):

i suppose that was a no

OpenStudy (anonymous):

Correct =)

OpenStudy (anonymous):

Completing the square is an interesting concept that is actually fairly easy to wrap your head around if you see it done geometrically.

OpenStudy (lgbasallote):

i'm not really used to doing completing square method

OpenStudy (anonymous):